OpenStudy (anonymous):

Whats (m4n2)1/2 square root of m2n2

OpenStudy (anonymous):

do you know this property \(\left(a^n\right)^{\frac{1}{m}}=a^{\frac{n}{m}}\)?

OpenStudy (anonymous):

\[\sqrt{m^4n^2}=\left(m^4\right)^\frac{1}{2}\left(n^2\right)^\frac{1}{2}=~?\]

OpenStudy (anonymous):

yeah I done that that's going to equal m2 n1... so do you times the m2 times m2 and n1 then times the n2?

OpenStudy (anonymous):

\[m2 n1*m2n1\]

OpenStudy (anonymous):

correct...

OpenStudy (anonymous):

thanks sooooo much!

OpenStudy (anonymous):

np... :)

OpenStudy (anonymous):

what about \[(8x3z9)2/3\]?

OpenStudy (anonymous):

almost the same rule, we just multiply the exponents... \[(a^m)^n=a^{m\times n}\]

OpenStudy (anonymous):

\[(8x^3z^9)^{\frac{2}{3}}=(8)^{\frac{2}{3}}(x^3)^{\frac{2}{3}}(z^9)^{\frac{2}{3}}=~?\]